OpenStudy (anonymous):

Help! How would I solve this 8x+12y=24?

OpenStudy (triciaal):

you cannot solve for 2 unknowns in one linear equation

OpenStudy (triciaal):

what is the original question?

OpenStudy (anonymous):

It says to graph 8x+12y=24 but you have to solve it first

OpenStudy (triciaal):

you can simplify by factoring out the common term 4 4(2x + 3y) = 4*6 2x + 3y = 6

OpenStudy (triciaal):